# Lambert’s problem in orbital dynamics: a self–contained introduction

### v1

@@ 3.3 Kepler’s laws of planetary motion
[…]
@@ 3.4 Brief historical digression
Kepler’s celebrated laws were published between 1609 and 1619 and had a tremendous impact on the astronomical community of the early $17^{th}$ century and beyond. They are the result of decades of careful measurements of the planet’s locations made by Tycho Brahe, and then of several more years of detailed mathematical analysis of the measurements (mostly of the orbit of Mars) by Kepler himself Wilson1972 . They are a testament of Brahe’s meticulous measurements (all of them made with the naked eye Tabak2011 , as the telescope which would not be around until 1609) as well as of Kepler’s mathematical and computational prowess and scientific integrity. Although by Kepler’s time the Heliocentric theory of the universe was starting to show serious cracks, the notion that planets followed circular orbits was still widely accepted as true. Kepler estimated the eccentricity of Mars’s orbit to be $e\approx 0.0926$ Xavier —which is not too far from that of a circle. Due to the coarseness of most contemporary measurements, Kepler could have easily attributed the discrepancy to measurement or computational error, however he trusted Brahe’s measurements and his own computations. Years of wrestling with the data had convinced him that the circle was not the correct answer and he pushed forward his elliptic discovery, even if it clearly went against the widely accepted theories.
@@
As originally stated by Kepler, they are purely kinematic descriptions that did not offer an underlying theory explaining the origin of the movement Hockey2014 ; Katz2008 . Due to the lack of a guiding principle leading Kepler’s analysis of empirical data, it can be argued that his conclusions were the product of serendipity: Brahe’s measurements were precise enough to show a discrepancy from a circular orbit, but not precise enough to reveal the perturbations due to the gravitational fields of other planets, which would have led Kepler astray Stillwell2010 . A complete dynamic explanation of planetary orbits would have to wait until Newton’s inverse square law, while the powerful and elegant analytical treatment that we have used above would require a further refinement of Newton’s calculus. The arguments and tools that we used in the previous section to obtain the equation of the orbits can be traced back to Euler and Lagrange in the mid to late $18^{th.}$ century Stillwell2010 .
@@
What we now call Kepler’s first law appeared in 1609, in his book Astronomia nova Kepler1609 , where he carefully describes the lengthy thought process that led him to the conclusion that Mars’s orbit is elliptic, and states that: {adjustwidth}3cm0cm “. . . the orbit of the planet is not a circle, but comes in gradually on both sides and returns again to the circle’s distance at perigee. They are all accustomed to call the shape of this sort of path ‘oval’.” Kepler1609 ; Linton2004 Kepler in fact derived an equation that is almost identical to the center–based equation for an ellipse, but using a different angular parameter that became known as the eccentric anomaly (depicted in Figure 5, where it is denoted as $\phi$) Linton2004 . After exhaustively working with the data for Mars and concluding that the ellipse was the only possible explanation, Kepler barely verified the data for other planets before generalizing the conclusion based on the fact that “the harmony of nature demanded that all have similar habitsBurton2010 . Regarding the location of the Sun at a focus of the ellipse, Linton Linton2004 points out that: “nowhere in the main body of the New Astronomy is the word ‘focus’ mentioned, and it was only later in his Epitome of Copernican Astronomy that Kepler emphasized this aspect of planetary orbits”. The Epitome of Copernican Astronomy, originally entitled “Epitome Astronomiae Copernicanae” Kepler1995 was intended to be a textbook in astronomy and was published in seven parts between 1618 and 1621 Caspar2012 ; Rothman2020 .
@@
As for the second law, Kepler derived it as an approximation, in an attempt to find a proxy for the time elapsed while a planet describes an arc. His derivation uses an argument that can be considered incipient integral calculus: as described by Katz Katz2008 , “Kepler then argued that the total time required to pass over a finite arc […] could be thought of as the sum of the radius vectors making up that part of the circle, or as the area swept by the radius vector”. Kepler was no stranger to “infinitesimal” arguments of that sort, which come remarkably close to integration. During his career he would use similar reasoning to approximate the area of a circle by considering the sum of infinitely many triangles with vertex on the center of the circle and sides along the circumference, as well as to calculate the surface area of a sphere by a similar argument involving infinitesimal cones Burton2010 . He famously went on to extend these techniques to calculating the volume of different solids of revolutions and wine barrels Kepler2018 .
@@
Going back to the equal areas law, Kepler discovered it sometime before 1605 and used it extensively in the calculations that would eventually lead to the first law, as he himself describes in his 1609 New Astronomy. However, his thoughts on this rule evolved over time and he eventually convinced that it was not an approximation, but a true statement. He would end providing a geometric proof that is almost correct by modern standards in 1621, as part of volume V of his Epitome of Copernican Astronomy Davis2003 .
@@ 4 Lambert’s problem
From the mathematical point of view, this question becomes a two–point boundary value problem involving the equations of motion (17) (or equivalently equation (10)). Boundary value problems differ from initial value problems in the sense that, instead of knowing the initial position and velocity, $\bm{r}$ and $\bm{\dot{r}}_{0}$, of the particle (or alternatively its energy and angular momentum), the initial and final positions $\bm{r}_{0}$ and $\bm{r}_{f}$ are prescribed. This condition changes substantially the mathematical treatment of the problem. One key difference is that boundary value problems tend to have multiple solutions even when the values at the boundaries are specified and therefore need supplemental physical information. Intuitively, a particle could travel along the same—closed—conic section either clockwise or counter–clockwise and complete multiple cycles before arriving at the final position. As we shall soon see, in our case the additional information will pertain to the number of revolutions that the particle can complete before arriving at the target endpoint. We will appeal to the geometric analysis of the possible trajectories that we developed in the previous section to sidestep the problem of integrating the boundary value problem.
In what follows, our analysis will be focused only on particles that follow elliptical trajectories. We will consider that the ellipse has a semi–major axis of length $a$, a semi–minor axis of length $b$, eccentricity $10\leq e<$ 1 (understanding the circle as the limiting case of ellipses as $e\to 0$), and the attracting center is located at the focus $F_{1}$ as depicted in Figure 5. Derivations following similar arguments for other conical orbits can be found in Grossman1996 ; Izzo2014 ; LaBl1969 ; LaBlDe1966 —although the exposition is less detailed.
Warning: As is clear from Figure 5 in this section the reference point has been shifted to the right focus. Therefore as pointed out in remark 1 every reference to equations (8) and (19a) should be understood as having the denominator $1+e\cos\theta$.
@@ 4.2 Kepler’s equation
ϕ˙=˙italic-$\dot{\phi}=\,$ $\frac{(1-e^{2})}{(1+e\cos\theta)^{2}}\cdot\frac{\dot{\theta}\sin\theta}{\sin\phi}\qquad\qquad$
@@
where the integration constant $-2\pi t_{0}/\mathcal{T}$ follows from the convention on setting the initial position at the perihelion (or periapsis), where the angle $\phi$ vanishes. This equation is known as Kepler’s equation.
The quantity $2\pi(t-t_{0})/\mathcal{T}$ appearing on the right hand side of (30) has angular units (recall that $2\pi=360^{\circ}$ is the angular measure of a circle in radians) and is called the mean anomaly. It is sometimes denoted by $M$ in the literature however, we will not follow that convention in the text. It can be interpreted as the angular position of a fictitious body that moves with constant angular velocity $2\pi/\mathcal{T}$ around a circular orbit of radius $a$. The fact that the mean anomaly is an angular measure that varies linearly with time enables the use of time intervals as proxies for angular displacement. If the actual angular displacement $\phi$ from the center of the ellipse is sought for, it can be obtained from the mean anomaly by solving Kepler’s equation. The angular displacement associated with the focus–centered polar coordinates $(r,\theta)$ can be obtained then from the eccentric anomaly from equation (28).
@@ 4.4 Lagrange’s solution
To connect the equality on the left of (35) and (36)(that involve only cosines) to equation (34c) (that involves only sines), we will make use of the trigonometric identity ${\displaystyle\cos\theta=1-2\sin^{2}(\theta/2)}$ which, upon substitution into said equations, yields
@@
where the positive and negative roots are determined by the difference $\theta_{2}-\theta_{1}$, as the cosine will yield positive and negative values depending on whether $\theta_{2}-\theta_{1}$ belongs to the interval $[0,\pi]$ or $[\pi,2\pi]$.
@@ 4.5 A second historical digression and some closing remarks
His interest in celestial mechanics dates back at least to 1744, when he got captivated by the great comet of Klinkenberg–Chéseaux DorregoLpez2023 which—by contemporary accounts—developed as many as six tails. It is unclear if the young Lambert obtained any mathematical results in the immediate aftermath of the comet’s passing, but we know that by early 1761 the idea behind Lambert’s theorem had matured in his mind enough for him to share it with Leonhard Euler in a letter dated on February 6th. 1762 Albouy2019 ; EulerArchive ; Bopp:1924 : {adjustwidth}3cm0cm “I forgot to turn problem §210 around, that the orbit may be found $1^{\circ}$ by the 3 sides $FN$, $FM$, $NM$ and the time $T$ required to traverse the arc $NM$. $2^{\circ}$ by the ratio $(FM:FN)$, the angle $NFM$, the time $T$ and the periodic time. If the diameter of the Sun can be measured precisely enough, two observations suffice to determine the Earth’s orbit using the latter theorem.” Very shortly after his letter to Euler, Lambert went on to treat the elliptic and hyperbolic cases and published the results as part of his book Lambert:1761 . He sent the book to Euler who, in March 24th. 1761 replied in a letter Albouy2019 ; EulerArchive ; Bopp:1924 : {adjustwidth}3cm0cm “The beautiful proof of the area of a parabolic sector, the expression for which you communicated to me gave me great pleasure; but I was even more surprised to see its application to elliptic sectors […] I easily recognize that the methods I proposed earlier may be improved considerably.” As indicated by the final phrase above, in fact Euler had started working on the determination of the movement of comets at least as far back as 1742, after the passing of Halley’s comet, publishing in 1743 a geometric method Euler1743 and then in 1744 the first purely analytic method to determine a parabolic orbit based on three observations of the comet Bistafa2021 ; Euler1744 . However, Euler did not attempt to generalize his calculations to orbits other than parabolic and therefore Lambert is generally credited with the introduction of the problem that now bears his name Albouy2019 .
@@
In his proof, Lambert made use solely of geometric arguments. This made the argument hard to understand and even harder to apply for practical calculations. In fact, in the same letter mentioned above, Euler points out that Albouy2019 ; EulerArchive ; Bopp:1924 : {adjustwidth}3cm0cm “Your theorem for expressing the area of a parabolic sector is excellent, I can see the truth of it, but by such detours, that I could never have arrived to it had I not known it in advance; I therefore wait impatiently to see the analysis leading to it without detours.” Lambert did not arrive at an alternate proof using the modern methods of calculus. This would have to wait until 1780, already after his passing, when Lagrange finally arrived at an argument similar to the one presented here. Countless other proofs to the theorem have been given and rediscovered over time. In the wonderful—but heavily mathematical—article Albouy2019 , Albouy discusses several “families” of proofs and presents a detailed timeline of early attempts and some highlights of modern ones.
@@
In his original statement, Lambert in fact did not pose a problem in the way we did at the beginning of Section 4). Instead, he made a statement about the relevant variables that determine the travel time, claiming that only the distances to the points, the length of the chord connecting them and the length of the semi–major axis of the conic (the parabolic case can be obtained by letting $a\to\infty$ (CoPr1993, , Problem 4.3)). In modern notation, the theorem can be stated as:
@@
Going back to the analysis in the previous section and recalling that the auxiliary variables $\alpha$ and $\beta$ are both functions of $r_{1}+r_{2}$, $c$, and the semi–major length $a$, it is clear then that equation (39) is only one step away from the form appearing in Theorem 4.1, which is why Lagrange’s argument constitutes an analytic proof of the statement.
Although clearly Theorem 4.1 is related to the orbit–determination Problem 1, they are certainly not the same thing. According to Albouy Albouy2019 , it was not until the 1960’s when the terms Lambert’s theorem and Lambert’s problem started being used interchangeably (with the second one eventually taking over). This was perhaps due to the fact that, with the advent of the space age, Lambert’s work was seen through the lens of orbital determination algorithms, many of which involved the solution of Lambert’s problem 1. As it concerns practical solution methods—rather than proofs of the theorem—there have been myriads of proposed algorithms. A recent comparative study of several of the most widely used approaches can be found in MG2021 .
Acknowledgments
All authors were partially funded by the U. S. National Science Foundation through the grant NSF-DMS-2137305. Lenox H. Baloglou thanks the generous support of the University of Arizona’s RII-Sponsored Campuswide Undergraduate Student–Initiated Original Research Program.

### v2

@@ 3.3 Kepler’s laws of planetary motion
$A=\;\tfrac{1}{2}\int_{\theta(t_{0})}^{\theta(t)}r^{2}d\theta\;=\;\tfrac{1}{2}\int_{t_{0}}^{t}r^{2}\dot{\theta}ds\underbrace{=}_{\text{\scriptsize Conservation}}\tfrac{1}{2}\int_{t_{0}}^{t}\tfrac{L}{m}\,ds\;=\;\tfrac{L}{2m}(t-t_{0}),$ (24)
@@ 3.4 Brief historical digression
Kepler’s celebrated laws were published between 1609 and 1619 and had a tremendous impact on the astronomical community of the early $17^{th}$ century and beyond. They are the result of decades of careful measurements of the planets’s locations made by Tycho Brahe, and then of several more years of detailed mathematical analysis of the measurements (mostly of the orbit of Mars) by Kepler himself Wilson1972 . They are a testament of Brahe’s meticulous measurements (all of them made with the naked eye Tabak2011 , as the telescope would not be around until 1609) as well as of Kepler’s mathematical and computational prowess and scientific integrity. Although by Kepler’s time the Heliocentric theory of the universe was starting to show serious cracks, the notion that planets followed circular orbits was still widely accepted as true. Kepler estimated the eccentricity of Mars’s orbit to be $e\approx 0.0926$ Xavier —which is not too far from that of a circle. Due to the coarseness of most contemporary measurements, Kepler could have easily attributed the discrepancy to measurement or computational error, however he trusted Brahe’s measurements and his own computations. Years of wrestling with the data had convinced him that the circle was not the correct answer and he pushed forward his elliptic discovery, even if it clearly went against the widely accepted theories.
@@
As originally stated by Kepler, his laws were purely kinematic descriptions that did not offer an underlying theory explaining the origin of the movement Hockey2014 ; Katz2008 . Due to the lack of a guiding principle leading Kepler’s analysis of empirical data, it can be argued that his conclusions were the product of serendipity: Brahe’s measurements were precise enough to show a discrepancy from a circular orbit, but not precise enough to reveal the perturbations due to the gravitational fields of other planets, which would have led Kepler astray Stillwell2010 . A complete dynamic explanation of planetary orbits would have to wait until Newton’s inverse square law, while the powerful and elegant analytical treatment that we have used above would require a further refinement of Newton’s calculus. The arguments and tools that we used in the previous section to obtain the equation of the orbits can be traced back to Euler and Lagrange in the mid to late $18^{th.}$ century Stillwell2010 .
@@
What we now call Kepler’s first law appeared in 1609, in his book Astronomia nova Kepler1609 , where he carefully describes the lengthy thought process that led him to the conclusion that Mars’s orbit is elliptic, and states that: “. . . the orbit of the planet is not a circle, but comes in gradually on both sides and returns again to the circle’s distance at perigee. They are all accustomed to call the shape of this sort of path ‘oval’.” Kepler1609 ; Linton2004 Kepler in fact derived an equation that is almost identical to the center–based equation for an ellipse, but using a different angular parameter that became known as the eccentric anomaly (depicted in Figure 5, where it is denoted as $\phi$) Linton2004 . After exhaustively working with the data for Mars and concluding that the ellipse was the only possible explanation, Kepler barely verified the data for other planets before generalizing the conclusion based on the conviction that “the harmony of nature demanded that all have similar habitsBurton2010 . Regarding the location of the Sun at a focus of the ellipse, Linton Linton2004 points out that: “nowhere in the main body of the New Astronomy is the word ‘focus’ mentioned, and it was only later in his Epitome of Copernican Astronomy that Kepler emphasized this aspect of planetary orbits”. The Epitome of Copernican Astronomy, originally entitled “Epitome Astronomiae Copernicanae” Kepler1995 was intended to be a textbook in astronomy and was published in seven parts between 1618 and 1621 Caspar2012 ; Rothman2020 .
@@
As for the second law, Kepler derived it as an approximation, in an attempt to find a proxy for the time elapsed while a planet describes an arc. His derivation uses an argument that can be considered incipient integral calculus: as described by Katz Katz2008 , “Kepler then argued that the total time required to pass over a finite arc […] could be thought of as the sum of the radius vectors making up that part of the circle, or as the area swept by the radius vector”. Kepler was no stranger to “infinitesimal” arguments of that sort, which come remarkably close to integration. During his career he would use similar reasoning to approximate the area of a circle by considering the sum of infinitely many triangles with vertex on the center of the circle and sides along the circumference, as well as to calculate the surface area of a sphere by a similar argument involving infinitesimal cones Burton2010 . He famously went on to extend these techniques to calculating the volume of different solids of revolution and wine barrels Kepler2018 .
@@
Going back to the equal areas law, Kepler discovered it sometime before 1605 and used it extensively in the calculations that would eventually lead to the first law, as he himself describes in his 1609 New Astronomy. However, his thoughts on this rule evolved over time and he eventually concluded that it was not an approximation, but a true statement. He would end providing a geometric proof that is almost correct by modern standards in 1621, as part of volume V of his Epitome of Copernican Astronomy Davis2003 .
@@ 4 Lambert’s problem
From the mathematical point of view, this question becomes a two–point boundary value problem involving the equations of motion (17) (or equivalently equation (10)). Boundary value problems differ from initial value problems in the sense that, instead of knowing the initial position and velocity, $\bm{r}$ and $\bm{\dot{r}}_{0}$, of the particle (or alternatively its energy and angular momentum), the initial and final positions $\bm{r}_{0}$ and $\bm{r}_{f}$ are prescribed. This condition changes substantially the mathematical treatment of the problem. One key difference is that boundary value problems tend to have multiple solutions and therefore need supplemental physical information to elucidate the correct answer. Intuitively, a particle could travel along the same—closed—conic section either clockwise or counter–clockwise and complete multiple cycles before arriving at the final position. As we shall soon see, in our case the additional information will pertain to the number of revolutions that the particle can complete before arriving at the target endpoint. We will appeal to the geometric analysis of the possible trajectories that we developed in the previous section to sidestep the problem of integrating the boundary value problem.
In what follows, our analysis will be focused only on particles that follow elliptical trajectories. We will consider that the ellipse has a semi–major axis of length $a$, a semi–minor axis of length $b$, eccentricity $0\leq e<1$ (understanding the circle as the limiting case of ellipses as $e\to 0$), and the attracting center is located at the focus $F_{1}$ as depicted in Figure 5. Derivations following similar arguments for other conical orbits can be found in Grossman1996 ; Izzo2014 ; LaBl1969 ; LaBlDe1966 —although the exposition in those sources is less detailed.
@@ 4.2 Kepler’s equation
$\dot{\phi}=\,$ $\frac{(1-e^{2})}{(1+e\cos\theta)^{2}}\cdot\frac{\dot{\theta}\sin\theta}{\sin\phi}\qquad\qquad$
@@
where the integration constant $2\pi t_{0}/\mathcal{T}$ follows from the convention on setting the initial position at the perihelion (or periapsis), where the angle $\phi$ vanishes. This equation is known as Kepler’s equation.
The quantity $2\pi(t-t_{0})/\mathcal{T}$ appearing on the right hand side of (30) has angular units (recall that $2\pi=360^{\circ}$ is the angular measure of a circle in radians) and is called the mean anomaly. It can be interpreted as the angular position of a fictitious body that moves with constant angular velocity $2\pi/\mathcal{T}$ around a circular orbit of radius $a$. It is sometimes denoted by $M$ in the literature. To avoid confusing it with the mass of the attractive center, we will not follow this convention in the text and will write out the full quotient explicitly instead. The fact that the mean anomaly is an angular measure that varies linearly with time enables the use of time intervals as proxies for angular displacement. If the actual angular displacement $\phi$ from the center of the ellipse is sought for, it can be obtained from the mean anomaly by solving Kepler’s equation. The angular displacement associated with the focus–centered polar coordinates $(r,\theta)$ can be obtained then from the eccentric anomaly from equation (28).
@@ 4.4 Lagrange’s solution
To connect the equality on the left of (35) and (36)(that involve only cosines) to equation (34c) (that involves only sines), we will make use of the trigonometric identity $\cos\theta=1-2\sin^{2}(\theta/2)$} which, upon substitution into said equations, yields
@@
where the positive and negative roots are determined by the difference $\theta_{2}-\theta_{1}$, as the cosine will yield positive and negative values depending on whether $\theta_{2}-\theta_{1}$ belongs to the interval $0,\pi]$ or $\pi,2\pi]$.
@@ 4.5 A second historical digression and some closing remarks
His interest in celestial mechanics dates back at least to 1744, when he got captivated by the great comet of Klinkenberg–Chéseaux DorregoLpez2023 which—by contemporary accounts—developed as many as six tails. It is unclear if the young Lambert obtained any mathematical results in the immediate aftermath of the comet’s passing, but we know that by early 1761 the idea behind Lambert’s theorem—under the assumption of a parabolic trajectory—had matured in his mind enough for him to share it with Leonhard Euler in a letter dated on February 6th. 1761: “I forgot to turn problem §210 around, that the orbit may be found $1^{\circ}$ by the 3 sides $FN$, FMFM, NMNM and the time $T$ required to traverse the arc $NM$. $2^{\circ}$ by the ratio $(FM:FN)$, the angle $NFM$, the time $T$ and the periodic time. If the diameter of the Sun can be measured precisely enough, two observations suffice to determine the Earth’s orbit using the latter theorem.” Albouy2019 ; Bopp:1924 ; EulerArchive Very shortly after his letter to Euler, Lambert went on to treat the elliptic and hyperbolic cases and published the results as part of his book Lambert:1761 . He sent the book to Euler who, in March 24th. 1761, replied in a letter: “The beautiful proof of the area of a parabolic sector, the expression for which you communicated to me gave me great pleasure; but I was even more surprised to see its application to elliptic sectors […] I easily recognize that the methods I proposed earlier may be improved considerably.” Albouy2019 ; Bopp:1924 ; EulerArchive As indicated by the final phrase above, in fact Euler had started working on the determination of the movement of comets at least as far back as 1742, after the passing of Halley’s comet, publishing in 1743 a geometric method Euler1743 and then in 1744 the first purely analytic method to determine a parabolic orbit based on three observations of the comet Bistafa2021 ; Euler1744 . However, Euler did not attempt to generalize his calculations to orbits other than parabolic and therefore Lambert is generally credited with the introduction of the problem that now bears his name Albouy2019 .
@@
In his proof, Lambert made use solely of geometric tools. This made the argument hard to understand and even harder to apply for practical calculations. In fact, in the same letter mentioned above, Euler points out that: “Your theorem for expressing the area of a parabolic sector is excellent, I can see the truth of it, but by such detours, that I could never have arrived to it had I not known it in advance; I therefore wait impatiently to see the analysis leading to it without detours.” Albouy2019 ; Bopp:1924 ; EulerArchive Lambert did not arrive at an alternate proof using the modern methods of calculus. This would have to wait until 1780, already after his passing, when Lagrange finally arrived at an argument similar to the one presented here. Countless other proofs to the theorem have been given and rediscovered over time. In the wonderful—but heavily mathematical—article Albouy2019 , Albouy discusses several “families” of proofs and presents a detailed timeline of early attempts and some highlights of modern ones.
@@
In his original statement, Lambert in fact did not pose a problem in the way we did at the beginning of Section 4. Instead, he made a statement about the relevant variables that determine the travel time, claiming that it depends only on the distances to the points, the length of the chord connecting them and the length of the semi–major axis of the conic (the parabolic case can be obtained by letting $a\to\infty$ (CoPr1993, , Problem 4.3)). In modern notation, the theorem can be stated as:
@@
Going back to the analysis in the previous section, and recalling that the auxiliary variables $\alpha$ and $\beta$ are both functions of $r_{1}+r_{2}$, $c$ and the semi–major length $a$, it is clear that equation (39) is only one step away from the form appearing in Theorem 4.1, which is why Lagrange’s argument constitutes an analytic proof of Lambert’s original statement.
Although clearly Theorem 4.1 is related to the orbit–determination Problem 1, they are certainly not the same thing. According to Albouy Albouy2019 , it was not until the 1960’s when the terms Lambert’s theorem and Lambert’s problem started being used interchangeably—with the second one eventually taking over. This was perhaps due to the fact that, with the advent of the space age, Lambert’s work was seen through the lens of orbital determination algorithms, many of which involved the solution of Lambert’s problem 1. As it concerns practical solution methods—rather than proofs of the theorem—there have been myriads of proposed algorithms. A recent comparative study of several of the most widely used approaches can be found in MG2021 .
5 Statements and declarations
Author contribution statment. Because of the nature of mathematical research, authors are considered on equal footing and are always listed in alphabetical order. In other words, as opposed to other disciplines, in mathematics there is no “first author”. Lenox Helene Baloglou, Parneet Gill and Tonatiuh Sánchez-Vizuet all contributed in equal proportion.
Competing interests. The authors declare that they have no conflict of interest.
Funding. All authors were partially funded by the United States National Science Foundation through the grant NSF-DMS-2137305. Lenox Helene Baloglou is thankful for the generous support of the University of Arizona’s RII-Sponsored Campuswide Undergraduate Student–Initiated Original Research Program.